An aqueous solution contains 0.483 M ammonia (NH3).

How many mL of 0.206 M hydroiodic acid would have to be added to 150 mL of this solution in order to prepare a buffer with a pH of 9.090?

mL
Answer #1

concentration of NH₃ = [NH₂] = 0.483 M = M, Volume of NH₃ = v= 150ml. milimoles of NH3 = M, V, = 0.483 x 150 = 72.45 mmol Con40 91 = 4.75 + log / 0.206 V2 (12:45- 0.2064 1.445= 0.2067₂ 72.45 - 012062 V = 207.85 ml Volume of HI should be added - (207.